I have tried this over 17 times and I still get the wrong answer when I check it.
4(12+z)-z=-192

Answer by algebrahouse.com(1659)

4(12 + z) - z = -192
48 + 4z - z = -192 {used distributive property}
48 + 3z = -192 {combined like terms}
3z = -240 {subtracted 48 from both sides}
z = -80 {divided both sides by 3}

Check:
4[12 + (-80)] - (-80) = -192
4(-68) + 80 = -192
-272 + 80 = -192
-192 = -192

For more help from me, visit: www.algebrahouse.com